OpenStudy (anonymous):

What is the surface area of a square pyramid with base edge 80 m and slant height 100 m?

OpenStudy (anonymous):

area of the square base of pyramid = 80 * 80 m^2 area of 4 triangles = 4((1/2)*80*100) m^2 Add them both and you get surface area of square pyramid